OpenStudy (anonymous):

Both trapezoids are similar. The area of the larger trapezoid is 61 ft2. Which is the best approximation for the area of the smaller trapezoid? 44 ft2 39 ft2 34 ft2 49 ft2

OpenStudy (anonymous):

I got 57 but that's not an answer choice... HELP!? Please..?

OpenStudy (anonymous):

Is there a picture that goes along with this question? If so, could you please add it?

OpenStudy (anonymous):

Ok. The two trapezoids are similar, which means the sides of the larger one are in a constant ratio to the sides of the smaller one. We divide 30/24 and find that the sides of the larger one are 1.25 times bigger than the smaller one. This should get you started.

OpenStudy (anonymous):

Okay.. So then i take it and multiply by 61 ft? \[\frac{ 24 }{ 30} = \frac{ x }{ 61 }\] \[24^{2}=841\] \[30^{2}=900\] So, then.. \[\frac{ 841 }{ 900 } = \frac{ x }{ 61 }\] \[900x=51,301\] \[\frac{ 51,301 }{ 900 }=57\] But.. that's not one of the answer choices..?